OpenStudy (august899):

Find the center and radius of the circle with the given equation. Then graph the circle. x^2+y^2=16?

hartnn (hartnn):

Circle with center = (0,0) and radius = r has the equation of the type: \(x^2 + y^2 = r^2 \)

hartnn (hartnn):

compare this standard equation with your question. r^2 = 16 so r = .. ?

OpenStudy (august899):

r=4?

hartnn (hartnn):

yes

OpenStudy (august899):

So do I do x^2+y^2=4

hartnn (hartnn):

you're asked about center and radius \(Center \equiv (0,0), r = 4 \)

OpenStudy (mathmale):

"So do I do x^2+y^2=4" Be a little more specific: "How do I transform x^2+y^2=4 into standard form and thus find the center and radius?"

OpenStudy (august899):

How did you find the center?

OpenStudy (mathmale):

Notice that both x and y are expressed as squares: x^2, y^2. This is equivalent to (x-0)^2, (y-0)^2.

hartnn (hartnn):

circle equation \(\Large (x-h)^2+ (y-k)^2 =r^2\) compare it with x^2+y^2 = r^2 h,k = 0,0 r = 4

OpenStudy (mathmale):

That's where the center, (0,0), came from.

OpenStudy (mathmale):

Borrowing from hartnn, \[\Large (x-h)^2+ (y-k)^2 =r^2\]

OpenStudy (mathmale):

contains the info you need about the center of the circle; it is the point (h,k).

OpenStudy (mathmale):

(x+3)^2 + (y-9)^2 = r^2 has center (-3,9). That's (h,k) in this situation.

OpenStudy (august899):

Okay. So the equation is (x-0)^2+(y-0)^2=4?

OpenStudy (mathmale):

that's partially correct. Your answer is equivalent to x^2 + y^2 = 2^2. Actually, you were given an equation with 16 on the right side. On the right side of this equation, you must type either 4^2 or 16. 4^2 is better. The radius is r=4, as previously discussed.

OpenStudy (august899):

Okay so (x-0)^2+(y-0)^2=16?